OpenStudy (anonymous):

MEDAL WILL BE GIVEN! Multiply the quantity of 2 square root 3 minus 1 times the quantity of 3 square root 3 minus 2

OpenStudy (queelius):

\[(2\sqrt{2}-1)(3\sqrt{3}-2)\] = \[(2)(3)\sqrt{2}\sqrt{3}-(2)\sqrt{2}(2)-(1)(3)\sqrt{3}+(1)(2)\] Can you take it from here?

OpenStudy (anonymous):

I sit okay for you to walk me through it? I mean f you can..?

OpenStudy (anonymous):

HELP?!

OpenStudy (queelius):

Sorry, I'm back. :)

OpenStudy (queelius):

We can walk through it one by one.

OpenStudy (queelius):

So, first, multiply the integers together, e.g., (2)(3) = 6.

OpenStudy (anonymous):

then what? @queelius

OpenStudy (queelius):

\[6\sqrt{6}-4\sqrt{2}-3\sqrt{3}+2\]

OpenStudy (anonymous):

im folloqing yah..continue

OpenStudy (queelius):

Well, I'm not sure much more can be productive done with it. Now, if you need an actual decimal representation, you're free to do that, e.g., this number is approximately equal to 5.8439317845.

OpenStudy (anonymous):

well these are the answer choices: negative 9 times the square root 3 16 plus 1 square root 3 13 times the square root 3 20 minus 7 square root 3

OpenStudy (queelius):

Well, that's not what we got at all. :) Could you repeat the initial problem? When I first responded to you, I wasn't sure if I correctly interpreted the question since there was some ambiguity in how to interpret the order of operations.

OpenStudy (anonymous):

you got a number wrong. Multiply the quantity of 2 square root 3 minus 1 times the quantity of 3 square root 3 minus 2

OpenStudy (queelius):

Ahh, yes, good catch.

OpenStudy (queelius):

That gives us 6*3-4*sqrt(3)-3*sqrt(3)+2

OpenStudy (queelius):

So, what else can we do to this? 6*3 = 18, and I also see two terms with sqrt(3) in it.

OpenStudy (queelius):

So, 18 - sqrt(3)(4 - 3) + 2 = ?

OpenStudy (anonymous):

don't you add 2 an 18?

OpenStudy (queelius):

Nope, we should add them together. What else can we do? sqrt(3) * (4 - 3) is equal to what?

OpenStudy (queelius):

Eerr, I meant, yup, we should add them together.

OpenStudy (anonymous):

4 sqr3 and -3 sqr3? and add them to get 4sqr3?

OpenStudy (queelius):

sqrt(3) * (4 - 3) = sqrt(3) * (1) = sqrt(3)

OpenStudy (queelius):

That gives us 20 + sqrt(3)

OpenStudy (queelius):

I'm not sure if that was one of the options. If not, we made a mistake somewhere. :)

OpenStudy (anonymous):

we kinda did its the last option I see 20 in it hahah

OpenStudy (queelius):

I see the error.

OpenStudy (queelius):

It should have been sqrt(3) * (-4 - 3) = -7sqrt(3)

OpenStudy (queelius):

So, our final answer is 20 - 7sqrt(3)

OpenStudy (queelius):

We have to be careful with our negative signs. :)

OpenStudy (anonymous):

thank you so much for the help!

OpenStudy (queelius):

Welcome.